OpenStudy (mjhpark25112001):

Choose the equation below that represents the line passing through the point (−5, 1) with a slope of three halves.

OpenStudy (mjhpark25112001):

can some one help

OpenStudy (mrnood):

do you know the 'slope intercept' form of the equation? y= mx +b ??

OpenStudy (mrnood):

'three halves' is just a slightly unusual way of writing 3/2

OpenStudy (mrnood):

hmm gone offline 2 minutes after posting....

OpenStudy (wanell):

Point slope y-y1=m(x-x1) where slope=m and you are given the point (x1,y1) slope=3/2 point is (-5,1) y-1=(3/2)(x+5) if you wan in slope intercept distribute y-1=3/2x+15/2 add 1 y=3/2x+17/2